OpenStudy (anonymous):

Find the indicated quantity, assuming u = 7i + j, v = i − 3j, and w = 3i + 5j. u · (v + w)

OpenStudy (lalaly):

\[v+w=4i+2j\]\[u.(v+w)=28+2=30\]

OpenStudy (anonymous):

thank you! :)

OpenStudy (lalaly):

:D